OpenStudy (bananas1234):

graph the function y = x^2 + 2X - 3 Draw a chart to show at least 3 ordered pairs that are on the graph. i also need to show all of my work

jimthompson5910 (jim_thompson5910):

Let's say x = 0. If x = 0, then what is the value of y?

OpenStudy (bananas1234):

0?

OpenStudy (p0sitr0n):

to graph this, you need to aknowledge that this is a parabola (2nd degree). it may also help to put it into cannonical form

jimthompson5910 (jim_thompson5910):

y = x^2 + 2x - 3 y = (0)^2 + 2(0) - 3 ... replace EVERY x with 0 y = ??

OpenStudy (bananas1234):

y = -3?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so we know that (0,-3) lies on this function

jimthompson5910 (jim_thompson5910):

when x = 1, y = ???

OpenStudy (bananas1234):

y = 0?